OpenStudy (anonymous):

In triangle XYZ, D and E are midpoints of XY and XZ. If YZ=1/2, find DE. b.what other relationships exists between DE and YZ? c.Name an angle congruent to angle XDE. Justify your answer.

OpenStudy (anonymous):



OpenStudy (anonymous):

So far, I have... a. DE=1.4 or 0.25 un. b. _______________________ c. angle XYZ because of corresponding angles postulate OR is it theorem??

Directrix (directrix):

The solution lies in the application of two Geometry theorems. ---- The segment joining the midpoints of two sides of a triangle is equal to half that of the third side and is parallel to the third side. a) DE = 1/2(1/2) = 1/4 b) Seg DE is parallel to Seg YZ If two parallel lines are cut by a transversal, corresponding angles are congruent. c) angle XDE is congruent to angle XYZ